# Boosting Transformer’s Robustness in PPG Signals Artifact Detection with Self-Supervised Learning

### v1

@@ V Conclusion
In the specialized context of PPG signal analysis for PICU applications, this investigation champions SSL as an innovative method to capitalize on the abundant yet underutilized unlabeled data. Through the adoption of SSL, the study has notably enhanced the Transformer model’s capacity to interpret signal characteristics, thereby improving its accuracy in the detection of PPG artifacts. The spectrum of SSL techniques explored—masking, contrastive learning, and DINO—revealed contrastive learning as the standout strategy, exhibiting remarkable efficacy, particularly with limited datasets.
The study further delves into the nuanced domain of contrastive learning, focusing on refining loss functions essential to SSL’s efficacy. Inspired by the foundational InfoNCE loss, a novel loss function has been devised to facilitate more stable training and superior convergence, subsequently boosting the Transformer’s performance in artifact identification tasks.
The research validates SSL’s potential to enhance neural network models’ performance, such as the Transformer, by effectively utilizing unlabeled data. This approach shows significant promise for deployment in environments like the PICU, where annotated data is a rarity. The study marks a critical evolution in machine learning, empirically validating the speculated advantages of SSL techniques over conventional supervised and unsupervised learning in enriching model performance and knowledge acquisition.

### v2

@@ V Conclusion
In the context of PPG signal analysis for PICU applications, this study presents SSL as a promising approach to utilizing the vast amount of unlabeled data. By integrating SSL into Transformer models, the results demonstrated improved capacity for signal interpretation and enhanced accuracy in detecting PPG artifacts. Among the various SSL techniques explored—masking, contrastive learning, and DINO—contrastive learning emerged as the most effective, particularly in scenarios with limited labeled data.
The research proposed refinements to contrastive learning to optimize SSL further, focusing on modifying the loss functions inspired by the original InfoNCE loss. These modifications enabled more stable training and better convergence, resulting in improved performance of the Transformer for artifact identification tasks.
However, this study is limited by its exclusive focus on PPG data collected within a single PICU environment, potentially restricting the generalizability of the findings. Future research should extend this work to more extensive and diverse benchmarking datasets [47, 24] to comprehensively evaluate the efficacy of SSL techniques in Transformer models for PPG analysis. Such validation will help establish the robustness and applicability of SSL approaches across varied clinical settings and datasets. Additionally, while our current approach using Transformer-based SSL achieves high accuracy in detecting PPG signal artifacts, it does not provide interpretability in its decision-making process. Integrating attention map analysis could help uncover which signal regions or patterns contribute most to artifact detection, thereby enhancing model transparency. Future work will explore this approach to improve interpretability, enabling a clearer understanding of the model’s decisions for clinical and practical applications.
This research underscores SSL’s transformative potential in elevating neural networks’ performance using limited labeled data, offering a viable path for future advancements in data-scarce environments like the PICU.